OpenStudy (anonymous):

Find the solution sets x^2+(x+1)^2=25 ?

OpenStudy (anonymous):

move 25 to the left hand side and expand the (x+1)^2 so you get x^2+x^2+2x+1-25=0 simplify it to get 2x^2+2x-24=0 take the 2 out to get 2(x^2+x-12)=0 factorise to get 2(x+4)(x-3)=0 so the solution are x=-4 and 3 factorise to get 2(x

OpenStudy (anonymous):

whoops sry, the last line was not intended haha.